OpenStudy (anonymous):

paula spots a glider located at an angle of elevation of 42 degrees. the distance between the glider and paula is 3280 feet. to the nearest foot, what is the height of the glider h from the ground?

OpenStudy (rulnick):

h / 3280 = sin(42)

OpenStudy (anonymous):



OpenStudy (rulnick):

h / 3280 = .66913 (approx) h = 3280 (.66913) = 2195 (approx) glider height is 2195 ft

OpenStudy (anonymous):

alright thanks. i was typing it wrong in my calculator.

OpenStudy (rulnick):

welcome